de = 22 + y dt = 2 + 2y 00 Paragraph B
Answer #1

Given question solved . Solne the system, 22 at ty dy = x+2y > 2x + y dy = x+ da dt => D2 =22ty 7 (0-2) x - y =o x+2y > Dy=x+2y 7 atzy - Dy = 0 => x-and if if f(D) x = 0 thers arxilary equation is -> (D-4D + 4-12 f (m) = am + bm tezo => (D? - 40+3)% =0 and if m=m), my are dthe soluation of the given system x(+) = gett eget y (k) -gette et anbitary vary arebitary where G and ey are ambil eans tant